OpenStudy (anonymous):

When does the relative max occur y= -x^3+ 3x^2-2 When does the relative max occur y= -x^3+ 3x^2-2 @Calculus1

OpenStudy (anonymous):

\[f(x)=-x^3+3x^2-2\] \[f'(x)=-6x^2+6x=-6x(x-1)\]

OpenStudy (anonymous):

critical points are ats x = 0, x = 1, and you can check that the first one is a relative min and the second is a relative max